OpenStudy (anonymous):

which of the numbers below is the sum of four consecutive whole numbers A) 15 B) 2000 C)2002 D)2004 E)2005

OpenStudy (anonymous):

NUMBER = x + (x+1) + (x+2) + (x+3) NUMBER = 4x + 6 NUMBER - 6 = 4x x = (1/4)(NUMBER - 6) x should be an integer (a whole number) for only one of the 5 options. the NUMBER of the 5 choices that you used to get an integer is your answer